OpenStudy (anonymous):

In 15 years, Maya will be twice as old as David is now. In 15 years, David will be as old as Maya will be 10 years from now. How old are Maya and David now?

OpenStudy (anonymous):

m+15 = 2d d+15= m + 10 d=20, m = 25

OpenStudy (anonymous):

wait how do i work the problem

OpenStudy (anonymous):

well the way i see it is that M+15 = 2d D+15M=10 tho

OpenStudy (anonymous):

so the way you would work it out is i belive add i am not sure hold on

OpenStudy (anonymous):

try what we did last time substitution m+15 = 2d so then m = 2d - 15 so in the second equation d + 15 = m + 10 use m= 2d -15 d+ 15 = (2d-15) + 10 you get d + 15 = 2d - 5 solve for d you get d = 20 then go back to m + 15 = 2d since d = 20 you get m + 15 = 2(20) so then m = 2(20) - 15 = 25

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

ok hold on he might be right but i belive there is another way an eser way

OpenStudy (anonymous):

there is another way but substitution is usually best if they cant get it...its a little longer but it can help them...you can use elimination and it would work just as well

OpenStudy (anonymous):

it works when they have trouble setting up or "seeing" the problem

OpenStudy (anonymous):

ik that but i am talking about anther way u have never did it then if so you would no

OpenStudy (anonymous):

dude its a system of equations......

OpenStudy (anonymous):

plus she is probably in algebra 1 or something so its elementary............you have to use methods that they can use..........you cant just pull out tricks from upper level math and expect them to reproduce the result.

OpenStudy (anonymous):

ok your right